# Generalization Bounds via Meta-Learned Model Representations: PAC-Bayes and Sample Compression Hypernetworks

### v2

@@ B.2 Second General PAC-Bayes Sample Compression Theorem
$\Delta\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q||P)+\ln\quantity(\frac{{\mathcal{A}}^{\prime}_{\Delta}(m)}{\delta})]$
@@
${\mathcal{A}}^{\prime}_{\Delta}(m)=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{(m-\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$
@@
We start from the end of the proof of Theorem B.1, just before Equation 10. As the loss of the hypothesis is computed on the data used to define $\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)$, this term cannot be bounded straightforwardly as before. Thus, we will assume that $\Delta$ is bounded to remove the spurious correlations. We want to remove the loss on $T_{\operatorname{\mathbf{i}}}$ in the following term.
@@
$=\frac{1}{m}\quantity[\operatorname{|\mathbf{j}|}\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega))+(m-\operatorname{|\mathbf{j}|})\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega))].$ (11)
@@
$\leq\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\frac{\eta\operatorname{|\mathbf{j}|}}{m}\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))+\frac{\eta(m-\operatorname{|\mathbf{j}|})}{m}\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$ (12)
@@
$=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{\eta\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\frac{\eta(m-\operatorname{|\mathbf{j}|})}{m}\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$
@@
We used Jensen’s inequality in Equation 12. Then, we have
${\mathcal{A}}_{\Delta}(m)=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{\eta\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$
Similarly to Theorem B.1, for all $\operatorname{\mathbf{j}}\in J$, we need to bound the moment generating function
To bound the moment generating function using the usual PAC-Bayes techniques, we need $\eta\leq m-\operatorname{|\mathbf{j}|}$ for all $\operatorname{\mathbf{j}}\in J$. Thus, the largest value of $\eta$ (which gives the tightest bound) that can be used is $\eta=m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}$. ∎
@@
$\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q||P)+\ln\quantity(\frac{\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}2\sqrt{m-\operatorname{|\mathbf{j}|}}}{\delta})]$
@@
$\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\leq\frac{\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))}{1-\frac{\lambda}{2}}+\frac{\mathrm{KL}(Q||P)+\ln\quantity(\frac{1}{\delta}\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}2\sqrt{m-\operatorname{|\mathbf{j}|}})}{\lambda(1-\frac{\lambda}{2})(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})}$
We can also recover a tighter bound than Theorem 39 of Germain et al. (2015). Instead of bounding the validation loss, they bound the training loss on the whole dataset, thus we can use Theorem B.2. Although this could theoretically lead to tighter bounds, later in the section we present a toy experiment that shows that in practice, it rarely is tighter.
@@
In the setting of Theorem B.1, with $\Delta(q,p)=2(q-p)^{2}$ and $S^{\prime}=S$, with probability at least $1-\delta$ over the draw $S\sim\operatorname{\mathcal{D}}^{m}$, we have
@@
This bound is tighter and more general than the results of Germain et al. (2015). Although their results claimed it held for all compression sets of size at most $\lambda$, only the case for compression set sizes $\lambda$ were considered. Indeed, in the proof of Lemma 38, with $\lambda>0$, when bounding $\mathcal{A}^{\prime}_{\Delta}(m)$, they consider the term
$\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\lambda}}\operatorname*{\mathbb{E}}_{S_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\lambda}}e^{(m-\lambda)2\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))-\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))^{2}}$ (13)
By the definition $S_{{\mathbf{j}}}=\{(\boldsymbol{x}_{j},y_{j})\}_{j\in{\mathbf{j}}}$, we have that $|S_{\operatorname{\mathbf{j}}}|=\operatorname{|\mathbf{j}|}$. Thus, Equation 13 only makes sense when $\lambda=\operatorname{|\mathbf{j}|}$. Moreover, later on, the decomposition of the loss
@@
$=\frac{1}{m}\quantity[\lambda\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))+(m-\lambda)\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))]$
If $\lambda\neq\operatorname{|\mathbf{j}|}$, then ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\neq\frac{1}{\lambda}\displaystyle\sum_{(\boldsymbol{x}_{i},y_{i})\in S_{\operatorname{\mathbf{j}}}}\ell(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)(\boldsymbol{x}_{i}),y_{i})$ and ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\neq\frac{1}{m-\lambda}\displaystyle\sum_{(\boldsymbol{x}_{i},y_{i})\in S_{\overline{\operatorname{\mathbf{j}}}}}\ell(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)(\boldsymbol{x}_{i}),y_{i})$. The result was used correctly in Germain et al. (2015), as they only considered this setting. However, the statement of the theorem was slightly incorrect.
@@
We now compare the bound on ^$\operatorname{\widehat{\mathcal{L}}}_{S}$ instead of ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}$. To do so, we compute the value of C.3 and a relaxed version of C.1 using Pinsker’s inequality. We choose $m=10000$, $\mathrm{KL}(Q||P)=100$, $\delta=0.01$, ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{j}}}}(h)=0$ and $J$ such that $\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}=2000$.
In this setting, we can investigate the relationship between the compression set size and the validation loss ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(h)$. In Figure 5, we report the difference between C.3 and the relaxed version of C.1.
Figure 5: This figure demonstrates the behavior of the difference between C.3 and the relaxed version of C.1, when the $\mathrm{KL}$ is small in comparison to the moment generating function. A positive value means C.3 is looser.
Figure 6: This figure demonstrates the behavior of the difference between C.3 and the relaxed version of C.1, when the $\mathrm{KL}$ is large in comparison to the moment generating function. A positive value means C.3 is looser.
It is obvious that in this setting, the exponential term $\exp\quantity(\frac{2(m-\operatorname{|\mathbf{j}|})\operatorname{|\mathbf{j}|}}{m})$ of C.3 is very penalizing. We can see that the difference becomes smaller when the validation loss becomes larger. This can be explained by the fact that the loss ^$\operatorname{\widehat{\mathcal{L}}}_{S}(h)=\frac{m-\operatorname{|\mathbf{j}|}}{m}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(h)$ decreases when $\operatorname{|\mathbf{j}|}$ becomes larger. However, it doesn’t come close to being advantageous. Indeed, it seems like C.3 only becomes advantageous when the $\mathrm{KL}$ is very large. In Figure 6, we chose $\mathrm{KL}(Q||P)=10000$. We then observe that for ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(h)$ larger than $0.6$, C.3 is actually smaller. However, the advantage doesn’t seem good enough to use this bound in any setting, as the advantage is only available in degenerate cases when the loss on the dataset is very high and the $\mathrm{KL}$ is also very large.
@@
Using Theorem B.1, we can obtain a new sample compression bound for real-valued messages. To do so, we restrict the result to use only Dirac measures as posterior distributions.

### v3

@@ B.2 Second General PAC-Bayes Sample Compression Theorem
$\Delta\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q||P)+\ln\quantity(\frac{{\mathcal{A}}^{\prime}_{\Delta}(m)}{\delta})],$
@@
${\mathcal{A}}^{\prime}_{\Delta}(m)=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{(m-\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}.$
@@
We start from the end of the proof of Theorem B.1, after using the i.i.d. assumption. As the loss of the hypothesis is computed on the data used to define $\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)$, this term cannot be bounded straightforwardly as before. To tackle this problem, we assume that $\Delta$ is bounded. We want to remove the need to compute the loss on examples $T_{\operatorname{\mathbf{j}}}=T\setminus T_{\overline{\operatorname{\mathbf{j}}}}$ in the following term.
$\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}\quad\mbox{(with $\eta>0$)}.$
@@
$=\frac{1}{m}\quantity[\operatorname{|\mathbf{j}|}\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega))+(m-\operatorname{|\mathbf{j}|})\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega))].$
@@
$\leq\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\frac{\eta\operatorname{|\mathbf{j}|}}{m}\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))+\frac{\eta(m-\operatorname{|\mathbf{j}|})}{m}\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$ (Jensen inequality)
@@
$=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{\eta\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\frac{\eta(m-\operatorname{|\mathbf{j}|})}{m}\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$ ($B\coloneqq\sup_{q,p\in[0,1]}\Delta(q,p)$)
@@
Similarly to Theorem B.1, we choose $\eta=m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}$ and we have
$\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{\eta\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$
$\leq\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{B\frac{(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\operatorname{|\mathbf{j}|}}{m}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{(m-\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}\eqqcolon{\mathcal{A}}^{\prime}_{\Delta}(m)\,.\qed$
@@
$\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q||P)+\ln\quantity(\frac{\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}2\sqrt{m-\operatorname{|\mathbf{j}|}}}{\delta})].$
@@
$\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\leq\frac{\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))}{1-\frac{\lambda}{2}}+\frac{\mathrm{KL}(Q||P)+\ln\quantity(\frac{1}{\delta}\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}2\sqrt{m-\operatorname{|\mathbf{j}|}})}{\lambda(1-\frac{\lambda}{2})(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})}.$
We can also recover a tighter bound than Theorem 39 of Germain et al. (2015). Instead of bounding the complement loss, they bound the training loss on the whole dataset, thus we can use Theorem B.2. Although this could theoretically lead to tighter bounds, later in the section we present a toy experiment that shows that in practice, it rarely is tighter.
@@
In the setting of Theorem B.2, with $\Delta(q,p)=2(q-p)^{2}$, with probability at least $1-\delta$ over the draw $S\sim\operatorname{\mathcal{D}}^{m}$, we have
@@
C.3 expresses a tighter and more general bound than the results of Germain et al. (2015). Moreover, as explained below, the latter is not always valid.
Identifying a small error in previous work of Germain et al. (2015).  Although the sample compression result of Germain et al. (2015) is stated for any compression sets of size at most $\lambda$, only the case for compression set sizes $\lambda$ were considered by the authors. Indeed, in the proof of Lemma 38, with $\lambda>0$, when bounding $\mathcal{A}^{\prime}_{\Delta}(m)$, they consider the term
$\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\lambda}}\operatorname*{\mathbb{E}}_{S_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\lambda}}e^{(m-\lambda)2\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))-\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))^{2}}.$ (8)
By the definition $S_{{\mathbf{j}}}=\{(\boldsymbol{x}_{j},y_{j})\}_{j\in{\mathbf{j}}}$, we have that $|S_{\operatorname{\mathbf{j}}}|=\operatorname{|\mathbf{j}|}$. Thus, Equation (missing) 8 only makes sense when $\lambda=\operatorname{|\mathbf{j}|}$. Moreover, later on, the decomposition of the loss
@@
$=\frac{1}{m}\quantity[\lambda\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))+(m-\lambda)\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))].$
If $\lambda\neq\operatorname{|\mathbf{j}|}$, then ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{j}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\neq\frac{1}{\lambda}\displaystyle\sum_{(\boldsymbol{x}_{i},y_{i})\in S_{\operatorname{\mathbf{j}}}}\ell(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)(\boldsymbol{x}_{i}),y_{i})$ and ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\neq\frac{1}{m-\lambda}\displaystyle\sum_{(\boldsymbol{x}_{i},y_{i})\in S_{\overline{\operatorname{\mathbf{j}}}}}\ell(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)(\boldsymbol{x}_{i}),y_{i})$.
The result was used correctly in Germain et al. (2015), as they only considered this setting. However, the statement of the theorem was slightly incorrect.
@@
Empirical comparison of the bounds.  We now compare the bound on ^$\operatorname{\widehat{\mathcal{L}}}_{S}$ instead of ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}$. To do so, we compute the value of C.3 and a relaxed version of C.1 using Pinsker’s inequality. We choose $m=10000$, $\mathrm{KL}(Q||P)=100$, $\delta=0.01$, ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{j}}}}(h)=0$ and $J$ such that $\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}=2000$. In this setting, we can investigate the relationship between the compression set size and the validation loss ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(h)$. In Figure LABEL:fig:pbsc_train_loss, we report the difference between C.3 and the relaxed version of C.1.
It is obvious that in this setting, the exponential term $\exp\!\big{(}\frac{2(m-\operatorname{|\mathbf{j}|})\operatorname{|\mathbf{j}|}}{m}\big{)}$ of C.3 is very penalizing. We can see that the difference becomes smaller when the validation loss becomes larger. This can be explained by the fact that the loss ^$\operatorname{\widehat{\mathcal{L}}}_{S}(h)=\frac{m-\operatorname{|\mathbf{j}|}}{m}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(h)$ decreases when $\operatorname{|\mathbf{j}|}$ becomes larger. However, it doesn’t come close to being advantageous. Indeed, it seems like C.3 only becomes advantageous when the $\mathrm{KL}$ is very large. In Figure LABEL:fig:pbsc_train_loss_large_kl, we chose $\mathrm{KL}(Q||P)=10000$. We then observe that for ^$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(h)$ larger than $0.6$, C.3 is actually smaller. However, the advantage doesn’t seem good enough to use this bound in any setting, as the advantage is only available in degenerate cases when the loss on the dataset is very high and the $\mathrm{KL}$ is also very large.
@@
Using Theorem B.1, we can obtain a new sample compression bound for real-valued messages. To do so, we restrict the result to use only Dirac measures as posterior distributions for the compression sets.